# LLM-enhanced Reranking in Recommender Systems

### v2

@@ Introduction
GMF (He et al. 2017) extends matrix factorization into a non-linear framework, serving as the primary global ranking method in this study. The GMF results represent recommendations prior to the application of any reranking process.
DLCM (Ai et al. 2018) enhances reranking efficacy by employing a recurrent neural network alongside an attention-based loss function to comprehend local ranking dynamics, aiming primarily at improving accuracy within recommendation outcomes.
PRM (Pei et al. 2019) leverages a transformer architecture with self-attention mechanisms to refine the entire recommendation list by acknowledging the inter-item influences, thereby concentrating on augmenting accuracy.
MMR (Carbonell and Goldstein 1998) aims to balance query relevance with the reduction of redundancy in reranked documents, employing a maximal marginal relevance score to bolster the diversity aspect in recommendation outcomes.
FastDPP (Chen, Zhang, and Zhou 2018) expedites Determinantal Point Processes (DPP) for Maximum A Posteriori (MAP) inference, facilitating the efficient production of diverse recommendation sets. This model focuses on the diversity aspect of the recommendation result.
FairRec (Wu et al. 2021) introduces a fairness-aware recommendation framework that employs decomposed adversarial learning and orthogonality regularization. It aims to alleviate bias concerning sensitive user attributes, thereby fostering fairness in recommendations without compromising overall performance.
RankGPT (Sun et al. 2023) investigates the application of LLMs in ranking tasks within information retrieval, employing a novel instructional permutation generation method alongside a sliding window strategy. This model is distinguished by its focus on accuracy. Note that, as a zero-shot LLM baseline, the permutation distillation method in the original paper is not implemented.
GoT (Besta et al. 2023) proposes a graph-of-thought approach to enhance LLMs’ prompting efficacy by structuring generated content as a graph. This facilitates synergistic outcomes, thought distillation, and feedback loop integration, aligning LLM reasoning more closely with human cognitive processes. Unlike LLM4Rerank, GoT adheres to predetermined node-to-node inference paths without historical data consideration. By applying a fixed path, “Accuracy-Diversity-Fairness-Stop.” GoT serves as a zero-shot LLM baseline focusing on the combination of accuracy, diversity, and fairness aspects in this paper.
Implementation Details
In the evaluation of the accuracy aspect, we adopt widely recognized metrics: Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG). For assessing the diversity aspect, we apply the commonly used metric -NDCG. To evaluate fairness, we use the Mean Absolute Difference (MAD) (Zhu, Hu, and Caverlee 2018). The MAD calculation is formalized as:
where and represent the predicted ratings for two distinct groups, and denotes the total number of ratings for group . Within the ML-1M dataset, we utilize the “genre” feature for diversity analysis and the “year” feature for fairness assessment, categorizing films into two groups based on their release year: pre-1996 and post-1996 (Zhu, Hu, and Caverlee 2018; Kamishima and Akaho 2017). For the KuaiRand dataset, “upload_type” serves as the criterion for diversity, while “video_duration”-divided into less than 60,000 ms and greater than 60,000 ms-serves for fairness evaluation. To fine-tune deep learning-based models for optimal performance, we set the learning rate at 0.001 and engage in a grid search to determine the best hyper-parameters. For zero-shot LLM baselines and LLM4Rerank, Llama-2-13B (Touvron et al. 2023) is selected as the default LLM backbone.
Overall Performance (RQ1)
@@
Aspect Combination Analysis (RQ2)
@@
Hyper-parameter Analysis
Recent studies have illuminated the challenges LLMs face in comprehensively processing long contexts laden with dense information (Liu et al. 2023). As the number of candidate items in a ranking sequence increases, so does the volume of semantic information, potentially overwhelming LLMs. This could explain the diminished efficacy observed when zero-shot LLMs are directly implemented in recommender systems that catalog millions of items. In light of this, in this section, we explore the impact of the hyper-parameter “candidate item number”, initially fixed at 20, on the reranking performance within the ML-1M dataset, as demonstrated in Figure 8 with LLM4Rerank-ADF.
The findings indicate a degradation in LLM4Rerank’s performance across various aspects as the “candidate item number” escalates. This outcome not only underscores the current limitations of LLMs in parsing long contexts but also reinforces their aptitude for tasks entailing more concise contextual information, such as reranking, over direct application in extensive recommendation or ranking frameworks. By constraining the number of items and focusing on rich semantic content within a single request, LLM4Rerank effectively narrows the semantic chasm across different aspect requirements, thus delivering more coherent reranking outcomes that enhance the overall recommendation quality.
Related Work
@@
Reranking in Recommendations
Reranking emerges as a critical post-processing strategy within recommender systems by introducing supplementary criteria to optimize the initial sequence of candidates. For example, the Deep Listwise Context Model (DLCM) (Ai et al. 2018) employs a recurrent neural network to sequentially process candidate items, thereby accruing contextual insights. Similarly, the Personalized Re-ranking Model (PRM) (Pei et al. 2019) utilizes the Transformer architecture (Vaswani et al. 2017) for encoding, enabling the modeling of extensive inter-item relationships. PRM enhances its efficacy by integrating personalized encoding mechanisms through the adoption of pre-trained embeddings.
Furthermore, the reranking phase is instrumental in addressing varied aspectual requirements, such as diversity (Yan et al. 2021; Gan et al. 2021) and fairness (Yadav, Du, and Joachims 2021; Raj and Ekstrand 2022). These aspects are increasingly recognized for their pivotal role in enriching user experience and ensuring that recommendations are congruent with overarching business objectives. For instance, the Maximal Marginal Relevance (MMR) (Carbonell and Goldstein 1998) model formulates a reranking schema emphasizing diversity by balancing document-query similarity against inter-document similarity. Conversely, FastDPP (Chen, Zhang, and Zhou 2018) innovates with an efficient greedy Maximum A Posteriori (MAP) inference for Determinantal Point Processes (DPP), facilitating the generation of both relevant and diverse recommendation sets. Additionally, FairRec (Wu et al. 2021) pioneers a fairness-focused recommendation framework employing decomposed adversarial learning and orthogonality regularization to mitigate biases and promote equity in the reranking process.
Moreover, recent advancements in LLMs indicate that LLM excels in reranking tasks with shorter contexts (Ma et al. 2023; Sun et al. 2023; Xiong et al. 2023; Lin et al. 2023) than in tasks with longer contexts. This discovery makes it possible to use LLM for better and more personalized reranking. For instance, DQ-LoRe (Xiong et al. 2023) uses dual queries for exemplar selection in prompting, RankGPT (Sun et al. 2023) employs a sliding window for ranking with long context, and the Graph of Thoughts (Besta et al. 2023) framework improves reranking outcomes with a fixed graph (Wei et al. 2022; Wang et al. 2022b) which allows for a complicated step-to-step data processing approach.
However, existing research typically focuses on one aspect, “Accuracy”, and occasionally includes one more aspect, such as “Diversity”. They overlook the need to integrate multiple aspects of different applications. In contrast, LLM4Rerank employs a function graph for automatic reranking, offering scalability and personalization in combining various aspects to suit different situations.
Conclusion

### v3

@@ 3.1.3. Implementation Details
3.2. Overall Performance (RQ1)
@@
3.3. Aspect Combination Analysis (RQ2)
@@ 3.5. Case Study
4. Related Work
@@
4.1. Reranking in Recommendations
Reranking emerges as a critical post-processing strategy within recommender systems (Zhao et al., 2018b, 2019; Liu et al., 2023c) by introducing supplementary criteria to optimize the initial sequence of candidates. For example, the Deep Listwise Context Model (DLCM) (Ai et al., 2018) employs a recurrent neural network to sequentially process candidate items, thereby accruing contextual insights. Similarly, the Personalized Re-ranking Model (PRM) (Pei et al., 2019) utilizes the Transformer architecture (Vaswani et al., 2017) for encoding, enabling the modeling of extensive inter-item relationships (Li et al., 2022; Liang et al., 2023).
Furthermore, the reranking phase is instrumental in addressing varied aspectual requirements, such as diversity (Yan et al., 2021; Gan et al., 2021) and fairness (Yadav et al., 2021; Raj and Ekstrand, 2022). These aspects are increasingly recognized for their pivotal role in enriching user experience and ensuring that recommendations are congruent with overarching business objectives. For instance, the Maximal Marginal Relevance (MMR) (Carbonell and Goldstein, 1998) model formulates a reranking schema emphasizing diversity by balancing document-query similarity against inter-document similarity. Conversely, FastDPP (Chen et al., 2018) innovates with an efficient greedy Maximum A Posteriori (MAP) inference for Determinantal Point Processes (DPP), facilitating the generation of both relevant and diverse recommendation sets. Additionally, FairRec (Wu et al., 2021) pioneers a fairness-focused recommendation framework employing decomposed adversarial learning and orthogonality regularization to mitigate biases and promote equity in the reranking process. Moreover, recent advancements in LLMs (Liu et al., 2024b; Xu et al., 2024a; Wang et al., 2023a) indicate that LLM excels in reranking tasks with shorter contexts (Ma et al., 2023; Sun et al., 2023; Xiong et al., 2023; Lin et al., 2023) than in tasks with longer contexts. This discovery makes it possible to use LLM for better and more personalized reranking. For instance, DQ-LoRe (Xiong et al., 2023) uses dual queries for exemplar selection in prompting, RankGPT (Sun et al., 2023) employs a sliding window for ranking with long context, and the Graph of Thoughts (Besta et al., 2023) framework improves reranking outcomes with a fixed graph (Wei et al., 2022; Wang et al., 2022a) which allows for a complicated step-to-step data processing approach.
However, existing research typically focuses on one aspect, “Accuracy”, and occasionally includes one more aspect, such as “Diversity”. They overlook the need to integrate multiple aspects of different applications. In contrast, LLM4Rerank employs a function graph for automatic reranking, offering scalability and personalization in combining various aspects to suit different situations. Further discussions about future developments are also provided in Appendix F.
5. Conclusion